# Fine-tuning Multimodal Transformers on Edge: A Parallel Split Learning Approach

### v1

@@ Abstract
Multimodal transformers integrate diverse data types like images, audio, and text, advancing tasks such as audio-visual understanding and image-text retrieval; yet their high parameterization limits deployment on resource-constrained edge devices. Split Learning (SL), which partitions models at a designated cut-layer to offload compute-intensive operations to the server, offers a promising approach for distributed training of multimodal transformers, though its application remains underexplored. We present MPSL [^1] Code will be made available upon acceptance., a parallel SL approach for computational efficient fine-tuning of multimodal transformers in a distributed manner, while eliminating label sharing, client synchronization, and per-client sub-model management. MPSL employs lightweight client-side tokenizers and a unified modality-agnostic encoder, allowing flexible adaptation to task-specific needs. Our evaluation across $7$ multimodal datasets demonstrates that MPSL matches or outperforms Federated Learning, reduces client-side computations by 250×250$\times$250 ×, and achieves superior scalability in communication cost with model growth. Through extensive analysis, we highlight task suitability, trade-offs, and scenarios where MPSL excels, inspiring further exploration.
@@ 1 Introduction
Figure 1: MPSL (Ours) vs. FL for multimodal learning on edge devices using Meta-Transformer, results averaged over 5 tasks.
@@ 2.1 Split Learning
Split Learning (SL) was introduced to overcome limitations of Federated Learning (FL), particularly the high computational demands on resource-constrained edge devices. By partitioning the model at a designated cut-layer, SL enables edge devices to process initial layers locally and send intermediate activations (smashed data) to a server for forward and backward propagation. Although this approach alleviates client-side computational burdens, SL requires sharing labels; impeding label privacy Gupta and Raskar (2018). To address this, variations like U-shaped SL were proposed to ensure that both input features and labels remain on the client-side by partitioning the model into three segments: client-head, server-body, and client-tail Gupta and Raskar (2018); Xu et al. (2023); Wang et al. (2023b). Despite these advancements, the sequential nature of vanilla and U-shaped SL imposed scalability limitations, prompting the development of Parallel Split Learning (PSL). PSL eliminates sequential dependencies by enabling multiple clients to train in parallel, significantly reducing training latency Jeon and Kim (2020); Baek et al. (2022); Oh et al. (2022); Chen et al. (2023); Han et al. (2021). Yet, early PSL works such as LocalSplitFed Han et al. (2021) require synchronization of client-side models – thereby introducing communication overhead – and often acquire parallelism by maintaining per-client sub-models on the server. The latter becomes infeasible when employing computationally intensive models with a large number of clients. However, subsequent PSL solutions acquire parallelism without this per-client sub-model constraint and introduce various optimizations for PSL, such as tackling the client decoupling problem identified in SGLR Pal et al. (2021) by using gradient aggregation techniques Pal et al. (2021); Oh et al. (2023); Lin et al. (2024), and exploring single-pass loss aggregation mechanisms Lyu et al. (2023) to minimize redundant computations and improve scalability. The advancements in PSL position the distributed learning framework as a viable alternative for the distributed training of computationally intensive models. Despite this, existing PSL frameworks primarily focus on unimodal tasks, leaving the integration of multimodal data largely unexplored. This gap highlights the need for advancements in PSL to effectively accommodate the unique challenges of training multimodal models in distributed settings.
2.2 Multimodal Architectures
@@
The emergence of multimodal transformers has revolutionized multimodal learning by facilitating the effective combination of diverse data modalities, including images, audio, and text. Here, models can be categorized into two main architectures: those using modality-specific encoders and those employing a single unified encoder. Models like CLIP Radford et al. (2021) use distinct encoders for each modality trained jointly with contrastive learning to align modality-specific embeddings into a shared space, demonstrating strong zero-shot transfer capabilities for tasks like image-text retrieval. Building on this, ImageBind Girdhar et al. (2023) supports multiple modalities by employing modality-specific tokenizers and encoders, generalizing across unseen modality pairs through pre-training on image-paired data. However, its reliance on separate encoders results in a large model size of $1.2$B parameters, limiting its suitability for resource-constrained environments Girdhar et al. (2023). Similarly, ViT-LENS Lei et al. (2023) aligns embeddings using lightweight modality-specific adapters with a frozen ViT encoder Dosovitskiy et al. (2020), reducing pre-training data requirements; yet retaining architectural complexity and size.
In contrast, unified encoder approaches like VATT Akbari et al. (2021) and Meta-Transformer Zhang et al. (2023) aim to share parameters across all supported modalities, enabling more scalable designs. VATT uses a modality-agnostic encoder with shared weights to process video, audio, and text, reducing parameter counts compared to modality-specific designs Akbari et al. (2021). Meta-Transformer extends this approach, employing a single encoder across twelve modalities using lightweight modality-specific tokenizers, showcasing promising results for parameter-efficient multimodal learning Zhang et al. (2023). Nonetheless, such models are predominantly designed for centralized training and remain computationally intensive for on-device learning.
@@
Distributed Multimodal Transformers. Distributed training of multimodal transformers has been explored in approaches like FedCLIP Lu et al. (2023), which uses lightweight adapters on top of a frozen backbone in an FL framework. However, the FL framework still requires clients to execute the frozen encoder backbone, limiting its applicability to resource-constrained edge devices. PSL offers a promising alternative; yet its application to multimodal transformers remains largely underexplored. Our work extends PSL to multimodal transformers, addressing client-side computational efficiency and modality alignment while overcoming prior PSL limitations like label sharing.
@@ 3.1 Notations
We use the following notations in the remainder of the paper. Let $\mathcal{N}=\{n_{1},...,n_{N}\}$ be the set of all clients, with $N$ as the total number of clients. For each client $n\in\mathcal{N}$, let $\mathcal{D}_{n}$ be its local dataset and $\mathcal{B}_{n}\subset\mathcal{D}_{n}$ a mini-batch of $\mathcal{D}_{n}$. The global batch is defined as $\mathcal{B}=\cup_{n\in\mathcal{N}}\mathcal{B}_{n}$. For a set of input modalities =$\mathcal{M}=\{m_{1},...,m_{M}\}$, let $T_{m}$ be a modality-specific tokenizer for each modality $m\in\mathcal{M}$. The multimodal transformer model, denoted as W, has weight parameters $\theta=(\theta_{1},\dots,\theta_{d})\in\mathbb{R}^{d}$ and is divided into three components: the head ($\textbf{W}_{h}$, modality-specific tokenizers), the body ($\textbf{W}_{b}$, unified encoder), and the tail ($\textbf{W}_{t}$, task-specific classifier). Hence, $\textbf{W}=[\textbf{W}_{h};\textbf{W}_{b};\textbf{W}_{t}]$, where $\textbf{W}_{h}$ consists of a set of tokenizers with $\textbf{W}_{h}=\{T_{m}\}_{m\in\mathcal{M}}$.
We split W into a server-side model $\mathcal{F}_{S}=[\textbf{W}_{b};\textbf{W}_{t}]$ and client-side model $\mathcal{F}_{C}=\textbf{W}_{h}$, with $\mathcal{F}_{C_{n}}$ as the client model for a client $n$. For a client $n$ with $M$ modalities, a single data sample is defined as $z=(\{x_{m_{i}}\}_{i}^{M},y)\in\mathcal{B}_{n}$, where $\{x_{m_{i}}\}_{i}^{M}=(x_{m_{1}},\dots,\textbf{x}_{m_{M}})$ represents the multimodal input, and $y$ denotes the label (i.e., the ground truth). The activations computed by $\mathcal{F}_{C_{n}}$ for a given input $\{x_{m_{i}}\}_{i}^{M}$ are denoted as $\textbf{a}_{n}$, while the predictions with respect to these activations computed by the server model $\mathcal{F}_{S}$ are expressed as $\hat{y}=\mathcal{F}_{S}(\textbf{a}_{n})$. Lastly, we define $\mathcal{L}_{C_{n}}$ as the $n$-th client’s loss function, and with $\mathcal{L}_{S}$ we denote the aggregated server-side loss function, given by $\mathcal{L}_{S}=\sum_{n\in\mathcal{N}}{\frac{|\mathcal{B}_{n}|}{|\mathcal{B}|}\mathcal{L}_{C_{n}}}$.
@@ 3.2 Modality fusion in MPSL
Processing multimodal inputs on clients. For a given client $n$ with a data sample $z=(\{x_{m_{i}}\}_{i}^{M},y)\in\mathcal{B}_{n}$, the client extracts the activations $\textbf{a}_{n}$ as follows:
$\textbf{a}_{n}=\mathcal{F}_{C_{n}}(\textbf{x}_{m_{1},m_{2}})=\begin{cases}[T_{m_{i}}(\textbf{x}_{m_{i}})]_{i}^{M}~{}~{}\text{if }\text{fusion}=early\\ {T_{m_{i}}(\textbf{x}_{m_{i}})}_{i}^{M}~{}~{}~{}~{}~{}~{}~{}\text{otherwise}\end{cases}$
where $T_{m_{i}}$ is the modality-specific tokenizer for modality $i$. The activations $\textbf{a}_{n}$ is subsequently send to the server, which in turn computes the prediction $\hat{y}$ with respect to $\textbf{a}_{n}$, and transmits back to the client. Given the label $y$ for $\textbf{x}_{m_{1},m_{2}}$, the client-side loss is computed as $\mathcal{L}_{C_{n}}(\hat{y},y)$ and send to the server to compute the gradients. Lastly, using the received cut-layer gradients from the server, client $n$ then performs backpropagation on its model $\mathcal{F}_{C_{n}}$. Formally, the client-side objective function is:
@@
Server-side training. We further elaborate on the training process on the server-side. Upon receiving the activations $\textbf{a}_{n}$ from a client $n$, the server computes its corresponding prediction $\hat{y}$ as follows:
[…]
The server computes the global aggregated loss $\mathcal{L}_{S}$ after receiving $\mathcal{L}_{C_{n}}$ from all clients $n\in\mathcal{N}$ and performs a single backward pass. This approach reduces server-side computational burden by requiring only one backward pass, regardless of number of clients, and eliminates the need to maintain multiple client sub-models. After backpropagation, the server sends the cut-layer gradients to all clients $n\in\mathcal{N}$. Formally, the server-side objective function is:
@@ 3.3 Post-training Model Construction
After training, the splitted multimodal transformer can be reassembled without restrictions. We impose no limitations in this regard; one approach is to construct a complete model for each client $n$ by concatenating the client-side model $\mathcal{F}_{C_{n}}$ with the server-side model $\mathcal{F}_{S}$, forming [$\mathcal{F}_{C_{n}};\mathcal{F}_{S}$]. This approach is suitable for personalization. Alternatively, a single client-side model $\mathcal{F}_{C_{\text{agg}}}$ can be obtained by aggregating all client-side models, following FedAvg McMahan et al. (2017), which is concatenated with the server-side model $\mathcal{F}_{S}$, resulting in [$\mathcal{F}_{C_{\text{agg}}}$;$\mathcal{F}_{S}$]. We adopt the latter approach to directly compare MPSL’s performance with other FL regimes McMahan et al. (2017); Lu et al. (2023). Notably, no client-side synchronization is enforced during the training phase.
@@ 4 Experiments
Models. Throughout our experiments, we use pre-trained weights from Meta-Transformer Zhang et al. (2023) based on ViT-B/16 Dosovitskiy et al. (2020), fine-tuning the complete encoder for image-text retrieval and the last six encoder blocks for classification tasks, unless stated otherwise. In Figure 6, where we explore various ViT encoder depths (Tiny, Small, Large, and Huge), we consistently fine-tune the latter half of the encoder blocks using pre-trained weights from Dosovitskiy et al. (2020). Late fusion was employed in all cases except for COCO-QA, T4SA, and MELD, where early fusion yields the best performance (see Table 5). For image-text retrieval we use the contrastive loss from Wang et al. (2023a), while for all classification tasks we employ Cross-Entropy loss.
@@
Baselines. We evaluate MPSL against centralized fine-tuning, standard FedAvg McMahan et al. (2017) and FedCLIP Lu et al. (2023), where lightweight modality-specific adapters are trained on a frozen backbone. The evaluation considers three criteria: task-specific performance (e.g., accuracy for classification and recall for retrieval tasks), client-side communication overhead (average MB transmitted per client in $1$ epoch), and computation overhead (measured in G-FLOPs per input sample and client-side trainable parameters). For rigorous evaluation, we run each experiment $3$ times, using a unique seed for each run, reporting the mean test accuracy.

### v2

@@ Abstract
Multimodal transformers integrate diverse data types like images, audio, and text, advancing tasks such as audio-visual understanding and image-text retrieval; yet their high parameterization limits deployment on resource-constrained edge devices. Split Learning (SL), which partitions models at a designated cut-layer to offload compute-intensive operations to the server, offers a promising approach for distributed training of multimodal transformers, though its application remains underexplored. We present MPSL [^1] Code is available on GitHub., a parallel SL approach for computational efficient fine-tuning of multimodal transformers in a distributed manner, while eliminating label sharing, client synchronization, and per-client sub-model management. MPSL employs lightweight client-side tokenizers and a unified modality-agnostic encoder, allowing flexible adaptation to task-specific needs. Our evaluation across $7$ multimodal datasets demonstrates that MPSL matches or outperforms Federated Learning, reduces client-side computations by 250×250$\times$250 ×, and achieves superior scalability in communication cost with model growth. Through extensive analysis, we highlight task suitability, trade-offs, and scenarios where MPSL excels, inspiring further exploration.
@@ 2 Related Work
Split Learning. Split Learning (SL) was introduced to address the limitations of FL, particularly the high computational demands on resource-constrained edge devices. By partitioning the model at a designated cut-layer, SL enables edge devices to process initial layers locally and send intermediate activations (smashed data) to a server for forward and backward propagation. Although this approach alleviates client-side computational burdens, SL requires sharing labels; impeding label privacy Gupta and Raskar (2018). To address this, variations like U-shaped SL were proposed to ensure that both input features and labels remain on the client-side by partitioning the model into three segments: client-head, server-body, and client-tail Gupta and Raskar (2018); Xu et al. (2023); Wang et al. (2023b). Despite these advancements, the sequential nature of vanilla and U-shaped SL imposed scalability limitations, prompting the development of Parallel Split Learning (PSL).
Early PSL methods like LocalSplitFed Han et al. (2021) required synchronizing client-side models, which added communication overhead, and maintaining a separate sub-model for each client on the server to enable parallelism – a setup that becomes impractical with many clients or large models. Subsequent PSL approaches overcome these limitations through several optimizations, such as using gradient aggregation Pal et al. (2021); Oh et al. (2023); Lin et al. (2024) to address client decoupling issues (as observed in SGLR Pal et al. (2021)), and employing single-pass loss aggregation Lyu et al. (2023) to reduce redundant computation and improve scalability. Such advancements position PSL as a viable alternative for the distributed training of computationally intensive models. Despite this, existing PSL frameworks primarily focus on unimodal tasks, leaving the integration of multimodal data largely unexplored. This gap underscores the pressing need to advance PSL toward effective multimodal learning, where multimodal challenges such as modality alignment are further complicated by the decentralized nature of distributed training.
@@
Multimodal Transformers. The emergence of multimodal transformers has enabled effective integration of diverse modalities such as images, audio, and text. Existing architectures typically follow two design paradigms: using modality-specific encoders or a unified encoder. Models like CLIP Radford et al. (2021) use separate encoders for each modality, aligning embeddings into a shared space via contrastive learning and demonstrating strong zero-shot transfer capabilities, particularly for image-text tasks. ImageBind Girdhar et al. (2023) extends this approach to additional modalities using modality-specific tokenizers and encoders, enabling generalization to unseen modality pairs through pretraining on image-paired data. ViT-LENS Lei et al. (2023) adopts a more lightweight strategy by attaching modality-specific adapters to a frozen ViT encoder Dosovitskiy et al. (2020), reducing data requirements while maintaining representational alignment. Despite their effectiveness, these models rely on modality-specific components or large backbone architectures, resulting in high parameter counts (e.g., up to $1.2$B for ImageBind) and architectural complexity – factors that limit their suitability for distributed or resource-constrained environments. Unified encoder approaches have emerged as a more scalable alternative, aiming to share parameters across modalities. VATT Akbari et al. (2021) employs a modality-agnostic encoder with shared weights to process video, audio, and text, reducing model size compared to modality-specific designs. Meta-Transformer Zhang et al. (2023) extends this idea by supporting $12$ modalities through a single encoder and lightweight modality-specific tokenizers, demonstrating strong performance in parameter-efficient multimodal learning. Nonetheless, such models are predominantly designed for centralized training and remain computationally intensive for on-device learning.
@@
Recently, parameter-efficient fine-tuning (PEFT) approaches have been explored for distributed training of multimodal transformers, with methods like FedCLIP Lu et al. (2023) applying lightweight adapters on top of a frozen backbone within a FL framework. However, FL still requires clients to execute the frozen encoder backbone, limiting its applicability to resource-constrained edge devices. PSL offers a promising alternative, yet its application to multimodal transformers remains largely underexplored. MPSL  extends PSL to multimodal transformers, addressing client-side computational efficiency and modality alignment, while also eliminating the need for label sharing.
@@ 3.1 Notations
We use the following notations in the remainder of the paper. Let $\mathcal{N}=\{n_{1},...,n_{N}\}$ be the set of all clients, with $N$ as the total number of clients. For each client $n\in\mathcal{N}$, let $\mathcal{D}_{n}$ be its local dataset and $\mathcal{B}_{n}\subset\mathcal{D}_{n}$ a mini-batch of $\mathcal{D}_{n}$. The global batch is defined as $\mathcal{B}=\cup_{n\in\mathcal{N}}\mathcal{B}_{n}$. For a set of input modalities =$\mathcal{M}=\{m_{1},...,m_{M}\}$, let $T_{m}$ be a modality-specific tokenizer for each modality $m\in\mathcal{M}$. The multimodal transformer model, denoted as $W$, has weight parameters $\theta=(\theta_{1},\dots,\theta_{d})\in\mathbb{R}^{d}$ and is divided into three components: the head ($W_{h}$, modality-specific tokenizers), the body ($W_{b}$, encoder), and the tail ($W_{t}$, task-specific classifier). Hence, $W=[W_{h};W_{b};W_{t}]$, where $W_{h}$ consists of a set of tokenizers with $W_{h}=\{T_{m}\}_{m\in\mathcal{M}}$.
We split $W$ into a server-side model $\mathcal{F}_{S}=[W_{b};W_{t}]$ and client-side model $\mathcal{F}_{C}=W_{h}$, with $\mathcal{F}_{C_{n}}$ as the client model for a client $n$. For a client $n$ with $M$ modalities, a single data sample is defined as $z=(\{\mathbf{x}_{m_{i}}\}_{i}^{M},y)\in\mathcal{B}_{n}$, where $\{\mathbf{x}_{m_{i}}\}_{i}^{M}=(\mathbf{x}_{m_{1}},\dots,\mathbf{x}_{m_{M}})$ represents the multimodal input, and $y$ denotes the label (i.e., the ground truth). The activations computed by $\mathcal{F}_{C_{n}}$ for a given input $\{\mathbf{x}_{m_{i}}\}_{i}^{M}$ are denoted as $\mathbf{a}_{n}$, while the predictions with respect to these activations computed by the server model $\mathcal{F}_{S}$ are expressed as $\hat{y}=\mathcal{F}_{S}(\mathbf{a}_{n})$. Lastly, we define $\mathcal{L}_{C_{n}}$ as the $n$-th client’s loss function, and with $\mathcal{L}_{S}$ we denote the aggregated server-side loss function, given by $\mathcal{L}_{S}=\sum_{n\in\mathcal{N}}{\frac{|\mathcal{B}_{n}|}{|\mathcal{B}|}\mathcal{L}_{C_{n}}}$.
@@ 3.2 Modality fusion in MPSL
Processing multimodal inputs on clients. For a given client $n$ with a data sample $z=(\{\mathbf{x}_{m_{i}}\}_{i}^{M},y)\in\mathcal{B}_{n}$, the client extracts the activations $\mathbf{a}_{n}$ as follows:
$\mathbf{a}_{n}=\mathcal{F}_{C_{n}}(\{\mathbf{x}_{m_{i}}\}_{i}^{M})=\begin{cases}[T_{m_{i}}(\mathbf{x}_{m_{i}})]_{i}^{M}~{}~{}\text{if }\text{fusion}=early\\ {T_{m_{i}}(\mathbf{x}_{m_{i}})}_{i}^{M}~{}~{}~{}~{}~{}~{}~{}\text{otherwise}\end{cases}$
where $T_{m_{i}}$ is the modality-specific tokenizer for modality $i$. The activations $\mathbf{a}_{n}$ are subsequently sent to the server, which in turn computes the prediction $\hat{y}$ with respect to $\mathbf{a}_{n}$, and transmits back to the client. Given the label $y$ for $\{\mathbf{x}_{m_{i}}\}_{i}^{M}$, the client-side loss is computed locally on the client-side as $\mathcal{L}_{C_{n}}(\hat{y},y)$ and sent to the server to compute the gradients. Since client $n$ only sends its loss to the server, label privacy is preserved. Lastly, using the received cut-layer gradients from the server, client $n$ then performs backpropagation on its model $\mathcal{F}_{C_{n}}$. Formally, the client-side objective function is:
@@
Server-side training. We further elaborate on the training process on the server-side. Upon receiving the activations $\mathbf{a}_{n}$ from a client $n$, the server computes its corresponding prediction $\hat{y}$ as follows:
$\hat{y}=\mathcal{F}_{S}(\mathbf{a}_{n})=\begin{cases}W_{t}(\mathrm{GAP}({W_{b}(\mathbf{a}_{n}))})~{}~{}\text{if }\text{fusion}=early\\ W_{t}(\mathrm{GAP}([W_{b}(\mathbf{a}_{n_{m_{i}}})]_{1}^{M}))~{}\text{otherwise}\end{cases}$
where $\mathrm{GAP}$ denotes the global average pooling operator. The server subsequently sends prediction $\hat{y}$ to client $n$. Note that in the case of late fusion, each modality in $\mathbf{a}_{n}$ is encoded independently using $W_{b}$, concatenated, and then processed with $\mathrm{GAP}$, ultimately leading to the final prediction $\hat{y}$ via $W_{t}$.
The server computes the global aggregated loss $\mathcal{L}_{S}$ after receiving $\mathcal{L}_{C_{n}}$ from all clients $n\in\mathcal{N}$ and performs a single backward pass on $\mathcal{F}_{S}$. This approach reduces server-side computational burden by requiring only one backward pass on the computationally intensive encoder $W_{b}$, regardless of number of clients, and eliminates the need to maintain multiple client sub-models. It is important to note that MPSL requires no explicit synchronization between client-side models. Instead, coordination emerges implicitly through shared cut-layer gradients computed from the aggregated global loss. This design mitigates client drift by aligning local updates around a common global signal and removes the need to manage individual client models on the server-side. After backpropagation, the server sends the cut-layer gradients to all clients $n\in\mathcal{N}$. Formally, the server-side objective function is:
@@ 3.3 Post-training Model Construction
After training, the splitted multimodal transformer can be reassembled without restrictions. We impose no limitations in this regard; one approach is to construct a complete model for each client $n$ by concatenating the client-side model $\mathcal{F}_{C_{n}}$ with the server-side model $\mathcal{F}_{S}$, forming [$\mathcal{F}_{C_{n}};\mathcal{F}_{S}$]. This approach is suitable for personalization. Alternatively, a single client-side model $\mathcal{F}_{C_{\text{agg}}}$ can be obtained by aggregating all client-side models, following FedAvg McMahan et al. (2017), which is concatenated with the server-side model $\mathcal{F}_{S}$, resulting in [$\mathcal{F}_{C_{\text{agg}}}$;$\mathcal{F}_{S}$]. We adopt the latter approach to directly compare MPSL’s performance with other FL regimes McMahan et al. (2017); Lu et al. (2023).
@@ 4 Experiments
Models. Throughout our experiments, we use pre-trained weights from Meta-Transformer Zhang et al. (2023) based on ViT-B/16 Dosovitskiy et al. (2020), fine-tuning the complete encoder for image-text retrieval and the last $6$ encoder blocks for classification tasks, unless stated otherwise. In Figure 6, where we use various ViT encoder depths (Tiny, Small, Large, and Huge), we fine-tune the latter half of the encoder blocks using pre-trained weights from Dosovitskiy et al. (2020). Late fusion was employed in all cases except for COCO-QA, T4SA, and MELD, where early fusion yields the best performance (see Table 5). For image-text retrieval we use the contrastive loss from Wang et al. (2023a), while for all classification tasks we employ Cross-Entropy loss.
@@
Baselines. We evaluate MPSL against centralized fine-tuning and two distributed learning baselines: standard FedAvgMcMahan et al. (2017) and FedCLIP Lu et al. (2023). As noted in Section 2, prior PSL work for distributed multimodal settings is scarce, motivating the development of MPSL. Since no existing PSL methods apply in this context, we focus our comparison on representative FL baselines: FedAvg, as a foundational method, and FedCLIP, which incorporates PEFT Houlsby et al. (2019) to enable efficient client-side adaptation. While other FL variants exist, they typically still depend on local model training followed by parameter aggregation and synchronization. We therefore compare MPSL to the most relevant and practical alternatives for distributed multimodal learning. We evaluate performance across $3$ criteria: (1) task-specific performance (e.g., accuracy for classification, recall for retrieval), (2) client-side communication overhead (average MB transmitted per client per epoch), and (3) computational overhead (measured in G-FLOPs per input and the number of trainable client-side parameters). All distributed baselines are trained for the same number of rounds to ensure a fair comparison. To reduce the effect of randomness during training, we report the mean test accuracy over $3$ runs with different random seeds.